# Finite-Precision Conjugate Gradient Method for Massive MIMO Detection

### v1

@@ V Simulation Results and Discussion
In this section, we will provide numerical results to verify our derived results. First, the simulation setup is provided. Then, we validate the insights from Remark 1 and 2 and the heuristic method of FP-CG detection. Finally, the superiority of FP-BJ-CG detection is verified.
@@ V-A1 Simulating Finite-Precision Arithmetic
The authors in [49] provided a function that can be utilized to simulate $\mathtt{fp32}$, $\mathtt{fp16}$, and other low-precision arithmetic. For the analysis of computational cost, we consider a ratio of $1:2:4$ for the costs of $\mathtt{fp16}$, $\mathtt{fp32}$, and $\mathtt{fp64}$ arithmetic [50].
@@ V-A2 Simulation Parameters
V-B Performance of FP-CG Detection
V-B1 Impact of Finite-Precision Arithmetic
As depicted in Fig. 1, we compare FP-CG detection in three cases: when only matrix-vector products use $\mathtt{bfloat16}$, only inner products use $\mathtt{bfloat16}$, and both use $\mathtt{bfloat16}$, under varying channel correlation in SNR $=20$ dB. The results indicate that using low-precision arithmetic for inner products does not impact the attainable accuracy of FP-CG detection, whereas matrix-vector products have a significant effect. Furthermore, as channel correlation increases, i.e., as the channel condition number grows, the gap and convergence delay between CG and FP-CG detection become more pronounced. These observations are consistent with the insights from Remark 1 and 2.
@@ V-B1 Impact of Finite-Precision Arithmetic
Figure 1: Convergence curve of FP-CG detection using $\mathtt{bfloat16}$ with different $\zeta$ in SNR $=20$ dB. MV: matrix-vector products using $\mathtt{bfloat16}$. IP: inner products using $\mathtt{bfloat16}$. MV + IP: both using $\mathtt{bfloat16}$.
V-B2 Heuristic Method
Fig. 2 illustrates the convergence curve of FP-CG detection employing the heuristic method for different correlation coefficients $\zeta$ in SNR $=20$ dB. For example, when $\zeta=0.5$, we first compute $u_{\rm MV}<1.33\times 10^{-3}$ using (21), then refer to Table I, and ultimately select $\mathtt{fp16}$. Compared to FP-CG detection with $\mathtt{bfloat16}$, the heuristic method ($\mathtt{fp16}$) achieves the same accuracy as CG detection. Moreover, Fig. 2 clearly shows that the heuristic method provides a practical guideline for selecting precision to ensure FP-CG detection convergence. Overall, the heuristic method serves as a valuable guideline for achieving attainable accuracy and ensuring convergence.
@@ V-B2 Heuristic Method
Figure 2: Convergence curve of FP-CG detection using the heuristic method with different $\zeta$ in SNR $=20$ dB.
@@ V-C1 Convergence Analysis
In Fig. 3, we compare the convergence curve of FP-BJ-CG Detection with different detection with $\zeta=0.8$ in SNR $=20$ dB, where SVD-CG detection applies the pre-processing method from [20] to CG detection. It can be observed that FP-BJ-CG detection achieves the fastest convergence among the compared methods. Since FP-BJ-CG detection has a lower condition number (See Theorem 5), $\mathtt{fp16}$ can be used for FP-BJ-CG detection based on the heuristic method instead of $\mathtt{fp32}$ for FP-CG detection, as shown in Fig. 2c.
Figure 3: Convergence curve of various detection using the heuristic method with $\zeta=0.8$ in SNR $=20$ dB.
@@ V-C2 BER Performance
As shown in Fig. 4, we compare the BER performance of FP-BJ-CG detection with different detection with $\zeta=0.8$ and $\mathcal{I}=10$. It can be observed that FP-BJ-CG detection with $\mathtt{fp16}$ exhibits similar performance to LMMSE detection at low SNR levels, with only almost $1.2$ dB performance loss at high SNR levels.
@@
where $\Delta{\bf H}\in\mathbb{C}^{M\times N}$ is the noise matrix. The ratio $\varepsilon$ between the power of channel and noise elements is set to be $20$ dB and $25$ dB. Since LMMSE detection under imperfect CSI achieves only sub-optimal performance, FP-BJ-CG detection also converges to a sub-optimal solution. Nevertheless, FP-BJ-CG detection achieves performance comparable to LMMSE detection under imperfect CSI, demonstrating its robustness.
(a) Various detection until convergence with $\zeta=0.8$ in SNR $=20$ dB, where FP-CG detection with $\mathtt{fp32}$ and FP-BJ-CG detection with $\mathtt{fp16}$.
Figure 6: Number of multiplications of various detection.
@@ V-C3 Computational Complexity Analysis
As demonstrated in Fig. 2c and 3, achieving the accuracy of LMMSE detection requires $15$, $17$, and $10$ iterations for CG, FP-CG, and FP-BJ-CG detection, respectively. Based on this, Fig. 6a compares the computational complexity of LMMSE, CG, FP-CG, and FP-BJ-CG detection until convergence under the same conditions as Fig. 3. Specifically, FP-CG detection with $\mathtt{fp32}$ reduces computational overhead by $43.3\%$ and $73.4\%$ compared to CG and LMMSE detection with $\mathtt{fp64}$, respectively. Similarly, FP-BJ-CG detection with $\mathtt{fp16}$ decreases $60.0\%$ and $81.3\%$ computational consumption compared to CG and LMMSE detection with $\mathtt{fp64}$, respectively.
Moreover, Fig. 6b shows the computational complexity of LMMSE, CG, FP-CG, and FP-BJ-CG detection with $10$ iterations, where FP-CG and FP-BJ-CG detection use $\mathtt{fp16}$. It can be observed that FP-BJ-CG detection with $\mathtt{fp16}$ decreases $40.0\%$ and $81.3\%$ computational cost compared to CG and LMMSE detection with $\mathtt{fp64}$, respectively. Although the computational complexity of FP-BJ-CG detection is a little higher than that of FP-CG detection, FP-BJ-CG detection achieves near BER performance than LMMSE detection as shown in Fig. 4 and 5.
@@ VI Conclusions
In this paper, we have proposed a low computational complexity CG detection from a finite-precision perspective. First, we have presented FP-CG detection to alleviate the computational bottleneck of each CG iteration and analyzed the corresponding performance to provide insights into the impact of finite-precision arithmetic. Then, a practical heuristic has been developed, which can choose suitable precisions to avoid significant degradation in FP-CG detection accuracy under different channel conditions. Furthermore, we have proposed FP-BJ-CG detection to simultaneously reduce the number of iterations and the computational complexity per iteration, and the corresponding analyses of convergence and computational costs have been provided. Finally, simulation results have demonstrated that FP-BJ-CG detection achieves similar BER performance to LMMSE detection at low SNR levels, with only almost $1.2$ dB performance loss at high SNR levels. Considering computation complexity, FP-BJ-CG detection decreases $40.0\%$ and $81.25\%$ of the computational cost compared to CG and LMMSE detection, respectively.
@@ Appendix A Proof of Theorem 2
Note that $01-\tau_{\rm W}>$ 0 for using $\mathtt{fp64}$, (47) can be expressed as
@@
When Algorithm 2 converges, we have [36]
@@ Appendix B Proof of Theorem 3
=$\overset{\left(a\right)}{=}\mathbb{E}\left(h_{mn}^{*}h_{mi}\right)\mathbb{E}\left(h_{on}h_{oi}^{*}\right)+\mathbb{E}\left(h_{mn}^{*}h_{on}\right)\mathbb{E}\left(h_{mi}h_{oi}^{*}\right)$
@@
where $(a)$ uses the Isserlis’ theorem for complex Gaussian variables [53, 54].
@@
=$\overset{\left(a\right)}{=}\sum_{m=1}^{M}{\mathbb{E}\left\{\left|h_{mn}\right|^{2}\right\}\mathbb{E}\left\{\left|h_{mi}\right|^{2}\right\}}+\frac{2}{M^{2}}\sum_{m=1}^{M-1}{\sum_{o=m+1}^{M}{\left[\mathbf{R}_{r}\right]_{mo}^{2}}}$
@@
(a) $n=i$ using $n=i=3$ as an example
(b) $n\neq i$ using $n=2$ and $i=4$ as an example
Figure 7: $\mathbb{V}\left\{\mathbf{h}_{n}^{\mathsf{H}}\mathbf{h}_{i}\right\}$ in (72) against BS antennas $M$ for $N=32$ and $\zeta=0.8$.
Further, we use simulation to verify the correctness of (72) as shown in Fig. 7. The analytical and simulated curves are very tight in all considered cases, confirming the correctness of our derived results.
Finally, from (72) and $10\leq\zeta<$ 1, when $M$ is asymptotically large, i.e., $M\longrightarrow\infty$, we obtain
@@
[…]
@@
Appendix C Proof of Theorem 5

### v2

@@ V Simulation Results and Discussion
In this section, we will provide numerical results to verify our derived results. First, the simulation setup is provided. Then, we validate the insights from Remarks 1 and 2 and the heuristic method of FP-CG detection. Finally, the superiority of FP-BJ-CG detection is verified.
@@ V-A1 Simulating Finite-Precision Arithmetic
The authors in [higham2019simulating] provided a function that can be utilized to simulate $\mathtt{fp32}$, $\mathtt{fp16}$, and other low-precision arithmetic. For the analysis of computational cost, we consider a ratio of $1:2:4$ for the costs of $\mathtt{fp16}$, $\mathtt{fp32}$, and $\mathtt{fp64}$ arithmetic [FAM].
@@ V-A2 Simulation Parameters
Figure 1: Convergence curve of FP-CG detection using $\mathtt{bfloat16}$ with different $\zeta$ in SNR =20=20 dB. MV: matrix-vector products using $\mathtt{bfloat16}$. IP: inner products using $\mathtt{bfloat16}$. MV + IP: both using $\mathtt{bfloat16}$.
@@
Figure 2: Convergence curve of FP-CG detection using the heuristic method with different $\zeta$ in SNR =20=20 dB.
V-B Performance of FP-CG Detection
V-B1 Impact of Finite-Precision Arithmetic
As depicted in Fig. 1, we compare FP-CG detection in three cases: when only matrix-vector products use $\mathtt{bfloat16}$, only inner products use $\mathtt{bfloat16}$, and both use $\mathtt{bfloat16}$, under varying channel correlation in SNR =20=20 dB. The results indicate that using low-precision arithmetic for inner products does not impact the attainable accuracy of FP-CG detection, whereas matrix-vector products have a significant effect on the attainable accuracy. Furthermore, as channel correlation increases, i.e., as the channel condition number grows, the gap and convergence delay between CG and FP-CG detection become more pronounced. These observations are consistent with the insights from Remarks 1 and 2.
V-B2 Heuristic Method
Fig. 2 illustrates the convergence curve of FP-CG detection employing the heuristic method for different correlation coefficients $\zeta$ in SNR =20=20 dB. For example, when $\zeta=0.5$, we first compute $u_{\rm MV}<1.33\times 10^{-3}$ using (21), then refer to Table I, and ultimately select $\mathtt{fp16}$. As shown in Fig. 2b, the heuristic method ($\mathtt{fp16}$) achieves the same accuracy as CG detection, while the use of $\mathtt{bfloat16}$ leads to a noticeable performance loss. Similar trends are observed in Figs. 2a and 2c. Moreover, Fig. 2 clearly shows that the heuristic method provides a practical guideline for selecting precision to ensure FP-CG detection convergence. Overall, the heuristic method serves as a valuable guideline for achieving attainable accuracy and ensuring convergence. Further, the simulation results confirm that the bound is tight, as performance degradation becomes evident when the precision does not hold the predicted threshold (the heuristic method).
@@ V-C1 Convergence Analysis
In Fig. 3, we compare the convergence curve of FP-BJ-CG Detection with different detection with $\zeta=0.8$ in SNR =20=20 dB, where SVD-CG detection applies the pre-processing method from [10653741] to CG detection. It can be observed that FP-BJ-CG detection achieves the fastest convergence among the compared methods. Since FP-BJ-CG detection has a lower condition number (See Theorem 5), $\mathtt{fp16}$ can be used for FP-BJ-CG detection based on the heuristic method instead of $\mathtt{fp32}$ for FP-CG detection, as shown in Fig. 2c.
Figure 3: Convergence curve of various detection using the heuristic method with $\zeta=0.8$ in SNR =20=20 dB.
@@ V-C2 BER Performance
As shown in Fig. 4, we compare the BER performance of FP-BJ-CG detection with different detection with $\zeta=0.8$ and $\mathcal{I}=10$. It can be observed that FP-BJ-CG detection with $\mathtt{fp16}$ exhibits similar performance to LMMSE detection at low SNR levels, with only an almost $1.2$ dB performance loss at high SNR levels.
@@
Figure 6: BER performance of various detection against SNR with convolutional coding, $\zeta=0.8$ and $\mathcal{I}=10$.
(a) Various detection until convergence with $\zeta=0.8$ in SNR =20=20 dB, where FP-CG detection with $\mathtt{fp32}$ and FP-BJ-CG detection with $\mathtt{fp16}$.
Figure 7: Number of multiplications for various detection.
@@
where $\Delta{\bf H}\in\mathbb{C}^{M\times N}$ is the noise matrix. The ratio $\varepsilon$ between the power of the channel and noise elements is set to be $20$ dB and $25$ dB. Since LMMSE detection under imperfect CSI achieves only sub-optimal performance, FP-BJ-CG detection also converges to a sub-optimal solution. Nevertheless, FP-BJ-CG detection achieves performance comparable to LMMSE detection under imperfect CSI, demonstrating its robustness.
In addition, Fig. 6 shows the BER performance of various detection methods against SNR under channel coding. We consider convolutional coding with a code rate of $\frac{1}{2}$ and employ a Viterbi decoder. The modulation scheme is 16-QAM. As illustrated in Fig. 6, the FP-BJ-CG detector using $\mathtt{fp16}$ precision achieves performance comparable to that of the LMMSE detector, demonstrating that the proposed FP-BJ-CG detection method performs effectively in coded massive MIMO systems.
@@ V-C3 Computational Complexity Analysis
As demonstrated in Figs. 2c and 3, achieving the accuracy of LMMSE detection requires $15$, $17$, and $10$ iterations for CG, FP-CG, and FP-BJ-CG detection, respectively. Based on this, Fig. 7a compares the computational complexity of LMMSE, CG, FP-CG, and FP-BJ-CG detection until convergence under the same conditions as Fig. 3. Specifically, FP-CG detection with $\mathtt{fp32}$ reduces computational overhead by $43.3\%$ and $73.4\%$ compared to CG and LMMSE detection with $\mathtt{fp64}$, respectively. Similarly, FP-BJ-CG detection with $\mathtt{fp16}$ decreases $53.3\%$ and $78.2\%$ computational consumption compared to CG and LMMSE detection with $\mathtt{fp64}$, respectively. Further, similar to the analysis in Example 1, the required computational cost for FP-BJ-CG detection is on the order of $2.7$ TFLOPS. Therefore, FP-BJ-CG detection requires only about $27.9\%$ of the computational resources of the NVIDIA A100, indicating that it can be efficiently implemented on a modern GPU.
Moreover, Fig. 7b shows the computational complexity of LMMSE, CG, FP-CG, and FP-BJ-CG detection with $10$ iterations, where FP-CG and FP-BJ-CG detection use $\mathtt{fp16}$. It can be observed that FP-BJ-CG detection with $\mathtt{fp16}$ decreases $30.0\%$ and $78.2\%$ computational cost compared to CG and LMMSE detection with $\mathtt{fp64}$, respectively. Although the computational complexity of FP-BJ-CG detection is a little higher than that of FP-CG detection, FP-BJ-CG detection achieves near BER performance than LMMSE detection, as shown in Figs. 4 and 5.
(a) Computational complexity of FP-CG and FP-BJ-CG as a function of $L$ with $N=32$ and $\zeta=0.8$ in SNR =20=20 dB.
(b) Computational complexity of FP-CG and FP-BJ-CG as a function of $N$ with $L=8$ and $\zeta=0.8$ in SNR =20=20 dB.
Figure 8: Trade-off between FP-CG and FP-BJ-CG detection.
Finally, Fig. 8 shows the simulated break-even points to further illustrate this trade-off between FP-CG and FP-BJ-CG detection. As depicted in Fig. 8a, for $N=32$, the break-even point occurs at $L=9$, indicating that FP-BJ-CG is more efficient for $L\leq 9$, while FP-CG performs better for $L>9$. Similarly, Fig. 8b shows that for $L=8$, the break-even point is $N=26$. Thus, for $N\geq 26$, FP-BJ-CG outperforms FP-CG, consistent with the theoretical insight from Remark 4.
@@ VI Conclusions
In this paper, we have proposed a low computational complexity CG detection from a finite-precision perspective. First, we have presented FP-CG detection to alleviate the computational bottleneck of each CG iteration and analyzed the corresponding performance to provide insights into the impact of finite-precision arithmetic. Then, a practical heuristic has been developed, which can choose suitable precisions to avoid significant degradation in FP-CG detection accuracy under different channel conditions. Furthermore, we have proposed FP-BJ-CG detection to simultaneously reduce the number of iterations and the computational complexity per iteration, and the corresponding analyses of convergence and computational costs have been provided. Finally, simulation results have demonstrated that FP-BJ-CG detection achieves similar BER performance to LMMSE detection at low SNR levels, with only almost $1.2$ dB performance loss at high SNR levels. Considering computation complexity, FP-BJ-CG detection decreases $30.0\%$ and $78.2\%$ of the computational cost compared to CG and LMMSE detection, respectively.
@@ Appendix A Proof of Theorem 2
Note that $1-\tau_{\rm W}>0$ for using $\mathtt{fp64}$, (52) can be expressed as
@@
When Algorithm 2 converges, we have [greenbaum1997estimating]
@@ Appendix B Proof of Theorem 3
$\overset{\left(a\right)}{=}\mathbb{E}\left(h_{mn}^{*}h_{mi}\right)\mathbb{E}\left(h_{on}h_{oi}^{*}\right)+\mathbb{E}\left(h_{mn}^{*}h_{on}\right)\mathbb{E}\left(h_{mi}h_{oi}^{*}\right)$
@@
where $(a)$ uses the Isserlis’ theorem for complex Gaussian variables [isserlis1918formula, 1057719].
@@
$\overset{\left(a\right)}{=}\sum_{m=1}^{M}{\mathbb{E}\left\{\left|h_{mn}\right|^{2}\right\}\mathbb{E}\left\{\left|h_{mi}\right|^{2}\right\}}+\frac{2}{M^{2}}\sum_{m=1}^{M-1}{\sum_{o=m+1}^{M}{\left[\mathbf{R}_{r}\right]_{mo}^{2}}}$
@@
Finally, from (77) and $0\leq\zeta<1$, when $M$ is asymptotically large, i.e., $M\longrightarrow\infty$, we obtain
@@
$\Longrightarrow$~ $\mathbf{H}^{\mathsf{H}}\mathbf{H}\xrightarrow{a.s.}\mathbf{R}_{t}.$ (82)
@@
Appendix C Comparison between BJ preconditioning and other preconditioning methods
In this appendix, we compare the advantages of the BJ preconditioner with both traditional preconditioning methods, such as the IC approach [7868573], and more recent methods, such as the UW-SVD preconditioner [10653741]. The comparison is conducted from two perspectives: convergence performance and computational complexity.
C-1 Convergence Performance
Fig. 9 compares the convergence behavior of BJ-, IC-, and UW-SVD-based CG detection under $\zeta=0.5$ at ${\rm SNR}=20$ dB. It can be observed that both BJ-CG and IC-CG achieve faster convergence than UW-SVD-CG detection.
Figure 9: Convergence curve of different preconditioning CG detection with $\zeta=0.5$ in SNR =20=20 dB.
C-2 Computational Complexity
Table II summarizes the computational complexity of generating different preconditioning matrices. The IC method exhibits an asymptotic complexity of $\mathcal{O}(N^{3})$, while the UW-SVD method scales linearly with the number of BS antennas $M$. In contrast, the BJ method requires substantially lower complexity, on the order of $\mathcal{O}(NL^{2})$, and is independent of $M$, thereby offering a favorable trade-off between performance and efficiency.
TABLE II: Computational Complexity of generating the different preconditioning matrices
Preconditioning methods BJ IC(1) UW-SVD
Computational Complexity $\mathcal{O}\left(NL^{2}\right)$ $\mathcal{O}\left(\left(N^{2}-2S\right)^{3/2}-\left(N^{2}-2S\right)^{1/2}\right)$ $\mathcal{O}\left(N_{\mathrm{UE}}MN\right)$
  • (1)
    $S$ is the number of zero entries.
Overall, the BJ preconditioner achieves competitive convergence performance with significantly lower computational complexity compared to both IC and UW-SVD methods, making it particularly well-suited for finite-precision massive MIMO detection.
Appendix D Proof of Theorem 5